# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v2

@@ 3.2 Regularization Induces Fractal Convergence Boundary
Finally, we present another implication of the chaos in gradient descent. In Appendix F, we show for the case $d=1$, i.e., $L(u,v)=\frac{1}{2}(uv-y)^{2}+\frac{\lambda}{2}(u^{2}+v^{2})$, $u,v\in\mathbb{R},\lambda\geq 0$, that, any continuous dynamical invariant must be constant. In particular, the imbalance $u^{2}-v^{2}$, which is known to be (approximately) preserved under gradient descent with small step sizes (Du et al.,, 2018; Arora et al.,, 2019), fails dramatically under large step sizes. Although this result does not directly extend to the case $d\geq 2$, we anticipate that the chaos strongly constrains the form of dynamical invariants.
4 General Matrix Factorization
We extend the results of Section 3 to matrix factorization (1). The key observation is that when the initialization lies in a particular slice $\mathcal{W}\subset\mathbb{R}^{2d\cdot d_{y}}$ of the parameter space, the trajectory stays in the slice and the dynamics decomposes into several parallel sub-dynamics in scalar factorization.
@@ 4 General Matrix Factorization
Consider gradient descent with step size $\eta$ for problem (1) with $d\geq d_{y}$. Consider the set $\mathcal{W}=\left\{(U,V)\in\mathbb{R}^{2d\cdot d_{y}}\colon\ \langle u^{i},u^{j}\rangle=\langle u^{i},v^{j}\rangle=\langle v^{i},v^{j}\rangle=0,\ \forall i\neq j\right\}$, where $u^{i},v^{i}$ denote the $i$th column of matrices $U,V$. The set $\mathcal{W}$ is forward-invariant, i.e., $\mathrm{GD}_{\eta}(\mathcal{W})\subset\mathcal{W}$. Moreover, if the initialization $(\bar{U},\bar{V})\in\mathcal{W}$, then for $i=1,\cdots,r$, the trajectory of the columns $(u^{i},v^{i})$ is identical to the trajectory of gradient descent applied to scalar factorization problem $L_{i}(u,v)=\frac{1}{2}(u^{\top}v-y_{i})^{2}+\frac{\lambda}{2}(\|u\|_{2}^{2}+\|v\|_{2}^{2})$, with step size $\eta$ and initialization $(\bar{u}^{i},\bar{v}^{i})$.
With this observation, all results presented in Section 3 extend verbatim to problem (1) with initializations in $\mathcal{W}$. We present the detailed extensions in Appendix G. Note, the slice $\mathcal{W}$ contains non-trivial, i.e., non-zero, initializations if and only if $d\geq d_{y}$. Thus, our results do not cover the low-rank setting. However, $\mathcal{W}$ contains the identity initializations $\bar{U}=\alpha I_{d},\bar{V}=\beta I_{d}$ for $d=d_{y}$, and the zero-asymmetric initialization (Wu et al.,, 2019). The identity initialization is common in training deep linear networks (Chou et al.,, 2024; Ghosh et al.,, 2025) and is closely related the training of linear residual networks (Hardt and Ma,, 2017; Bartlett et al.,, 2018). Our results therefore also apply in those setting. Although our results are established for initializations in $\mathcal{W}$, we experimentally verify that the chaotic phenomena persist over general initializations and in deep matrix factorization; details are provided in Appendix G.
Discussion of the mechanism
We provide an intuitive explanation for the emergence of chaos in matrix factorization. Although the update map $\mathrm{GD}_{\eta}$ is non-invertible, in Appendix G we show that the parameter space is partitioned by a measure-zero set $\mathcal{K}_{\eta}$, so that the restriction of $\mathrm{GD}_{\eta}$ to each connected component of $\mathbb{R}^{2d\cdot d_{r}}\setminus\mathcal{K}_{\eta}$ is a covering map. This measure-zero set comprises primarily the critical points of the map $\mathrm{GD}_{\eta}$, $\left\{\theta\in\mathbb{R}^{2d\cdot d_{y}}\colon\det J(\mathrm{GD}_{\eta}(\theta))=\det(I-\eta\nabla^{2}L(\theta))=0\right\}$, which are the points where the Hessian of the loss has an eigenvalue equal to $1/\eta$. We reason that chaotic phenomena arise when there exists a connected component $\mathcal{C}\subset\mathbb{R}^{2d\cdot d_{y}}\setminus\mathcal{K}_{\eta}$ that satisfies (i) $\mathcal{C}$ contains at least two global minimizers, and (ii) $\mathcal{C}\subseteq\mathrm{GD}_{\eta}(\mathcal{C})$. To see this, assume that $\mathrm{GD}_{\eta}$ has covering degree $m$. For any minimizer $p\in\mathcal{C}$, since $p\in\mathcal{C}\subseteq\mathrm{GD}_{\eta}(\mathcal{C})$, there exists a neighborhood $O_{p}$ of $p$ such that $\mathrm{GD}^{-1}(O_{p})$ has $m$ connected components that are contained in $\mathcal{C}$. By induction, the preimage of $O_{p}$ under the $N$th iteration map, $\mathrm{GD}^{-N}(O_{p})$, has $m^{N}$ components. Then the basin of attraction of points near $p$ is $B(p)=\cup_{N=1}^{\infty}\mathrm{GD}^{-N}(O_{p})$, which has infinitely many components. These accumulate at the convergence boundary, as the boundary often contains an invariant set with expanding directions and thus attracts inverse dynamics. Since the same phenomenon holds for all global minimizers in $\mathcal{C}$, their associated basins tend to be intricately interwoven near the convergence boundary, which gives rise to sensitive dependence on the initialization (see right panel of Figure 1). Additionally, by definition, $B(p)=\mathrm{GD}_{\eta}(B(p))$, which means that $B(p)$ can be folded $m$ times onto itself. This characterizes a self-similar structure. We point out that although this heuristic provides an intuitive explanation for the emergence of chaos, a rigorous investigation is left for future work.
@@ 5 Conclusion
We offered a rigorous characterization of gradient descent with large step sizes in matrix factorization. Our results reveal two striking phenomena: near the convergence boundary, the selection of the minimizer is unpredictable, and adding regularization can induce a fractal convergence boundary that makes the convergence itself unpredictable. As a driver of this complexity, we suggested a covering map structure exhibited by the gradient update map on parameter regions where the inverse step size is not an eigenvalue of the Hessian of the loss.
@@
Although our characterizations substantially expand the state of knowledge in non-convex overparametrized optimization in the particular setting of matrix factorization, further research is needed to rigorously characterize the dynamics of large-step-size gradient descent in other settings, such as general initializations, deep matrix factorization, or neural networks with nonlinear activation functions. We believe the contributed insights can aid in the development of such program.
@@
Code to reproduce our experiments is made available at https://anonymous.4open.science/r/chaos-matrix-factorization-07C5.
@@ Acknowledgment
This project has been supported by NSF CCF-2212520 and NSF DMS-2145630. GM has also been supported by DARPA in AIQ project HR00112520014, DFG in SPP 2298 (FoDL) project 464109215, and BMFTR in DAAD project 57616814 (SECAI). We thank Ryan Tibshirani for insightful discussions.
@@
  • Appendix I: Additional experiments
@@ A.1 Chaotic dynamical systems
A common definition of chaotic dynamical systems is by Devaney and Eckmann, (1987). A dynamical system $F\colon X\to X$, where $X$ is the state space, is chaotic if: (i) it is sensitive to initialization, (ii) it is topological transitive, and (iii) periodic points are dense in $X$. Here, sensitivity to initialization requires that there exists $\delta>0$ such that for any $x\in X$ and any neighborhood of $x$, there exists $N>0$ and $y$ in the neighborhood such that $\operatorname{dist}(F^{N}(x),F^{N}(y))>\delta$. This is weaker than the property we presented in Theorem 1 and Theorem 3: when the converged points of two trajectories are different, the trajectories must differ by a positive difference at some time $N$, but not vice versa. In Proposition 14, we show that the boundary $\partial\mathcal{D}_{\eta}^{\prime}$, as defined in Theorem 1, is invariant under gradient descent. In Proposition 18, we show that when restricted to $\partial\mathcal{D}_{\eta}^{\prime}$, the gradient descent system is semi-conjugate to a one-dimensional system that is precisely Devaney chaotic. However, we show in Proposition 19 that the original gradient descent system is not Devaney chaotic when $d\geq 2$ as it fails to be topological transitive.
A system $F\colon X\to X$ is topological transitive if for any pair of non-empty open sets $U,V$, there exists $N$ such that $F^{N}(U)\cap V\neq\varnothing$. A family of dynamical systems that exhibit transitivity is the family of Axiom-A diffeomorphisms (Smale,, 1967). These are dynamical systems where the set of non-wandering points is hyperbolic and is equal to the closure of the set of periodic points. A closed set $\Lambda$ is hyperbolic if it is forward invariant and at each point $x\in\Lambda$ the tangent space of the ambient space splits as a direct sum of stable and unstable subspaces. It is known that an Axiom-A diffeomorphism is always transitive on each of its basic set (Bowen et al.,, 2008, Chapter 3). However, gradient update maps typically are not expected to satisfy this definition as they in general are not global diffeomorphisms.
Another definition of chaotic dynamical system is by Li and Yorke, (1975). They considered a dynamical system $F\colon X\to X$ with $X\subset\mathbb{R}$ being an interval chaotic if $F$ has a periodic orbit with period three. They showed that if such a periodic orbit exists, then (i) $F$ has periodic orbit with any period; (ii) there exists an uncountable set $S\subset J$ such that, for every $p,q\in S$ with $p\neq q$,
@@
In Proposition 17, we showed that the restricted system $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}}$ is semi-conjugate to a one-dimensional system that is Li-Yorke chaotic. In general, Devaney chaos and Li-Yorke chaos do not imply each other. For a detailed comparison between different notions of chaotic dynamical systems, see Elaydi, (2007).
@@ A.2 Fractal basin boundary
The fractal convergence boundary studied in this work falls into a more general notion, called a fractal basin boundary. The seminal classification given by McDonald et al., (1985) divides fractal basin boundaries into three categories: quasicircles, locally connected but not quasicircles, and locally disconnected. The most regular type, quasicircle, is typical in the Julia sets of complex analytic maps. However, as noted by McDonald et al., (1985), properties of complex analytic maps do not generalize to real maps, and hence quasicircle is uncommon in real systems. We observe that the convergence boundary in our study falls in the second category, locally connected but not quasicircles. This type of boundary has been observed in several planar maps, i.e., dynamical systems defined on regions of $\mathbb{R}^{2}$. A well-known example is the following system:
@@
The basin boundary of this system is precisely the Weierstrass curve. McDonald et al., (1985) argued that a typical characteristic of this type of boundaries is the local stratification structure, which also appears in the convergence region in our case (see left panel in Figure 1). To our knowledge, however, all examples of locally connected boundaries appearing in the literature, including those presented by McDonald et al., (1985); Hunt et al., (1999); Rosa Jr and Ott, (1999), are bounded, whereas the convergence region in our case is shown to be unbounded. Classical approaches do not apply to our study, as most of those theoretical studies are case-specific. The last category has the most complicated structure and, as noted by Aguirre et al., (2009), turns out to appear more commonly in physical systems. Boundaries in this category typically exhibit a Cantor set structure. Examples include the famous Hénon map and the horseshoe map. For a recent review of the fractal boundaries, we refer readers to Aguirre et al., (2009).
Appendix B Measure of chaos in dynamical system

### v3

@@ 3.2 Regularization Induces Fractal Convergence Boundary
Finally, we present another implication of the chaos in gradient descent. In Appendix F, we show for the case $d=1$, i.e., $L(u,v)=\frac{1}{2}(uv-y)^{2}+\frac{\lambda}{2}(u^{2}+v^{2})$, $u,v\in\mathbb{R},\lambda\geq 0$, that, any continuous dynamical invariant must be constant. In particular, the imbalance $u^{2}-v^{2}$, which is known to be (approximately) preserved under gradient descent with small step sizes (Du et al., 2018; Arora et al., 2019), fails dramatically under large step sizes. Although this result does not directly extend to the case $d\geq 2$, we anticipate that the chaos strongly constrains the form of dynamical invariants.
4 Matrix Factorization and Beyond
In this section, we first extend results in Section 3 to general matrix factorization with orthogonal initializations. We then analyze the underlying mechanism that gives rise to the chaotic phenomena, and discuss to what extent this mechanism may extend to more general settings. Finally, we present experiments showing chaos in a real-world machine learning setting.
4.1 Matrix Factorization with Orthogonal Initializations
Consider gradient descent in matrix factorization (1) with initializations in the following subspace:
where $u^{i},v^{i}$ denote the $i$th column of $U,V$. The subspace $\mathcal{W}$ includes several commonly studied initialization schemes, such as the scaled identity initializations $\bar{U}=\alpha I_{d},\bar{V}=\beta I_{d}$ (Chou et al., 2024; Ghosh et al., 2025), zero-asymmetric initialization (Wu et al., 2019), and those used in training linear residual networks (Hardt and Ma, 2017; Bartlett et al., 2018). Note that in the low-rank setting $d<d_{y}$, constraints in (5) require at least $d_{y}-d$ pairs of $(u^{i},v^{i})$ to be initialized to zero.
A key observation is that results of Section 3 precisely characterize the gradient descent dynamics on $\mathcal{W}$: with initialization in $\mathcal{W}$, the trajectory remains in $\mathcal{W}$, and the dynamics decouple column-wise: each pair $(u^{i},v^{i})$ evolves independently according to the scalar factorization dynamics with target $y_{i}$, the $i$th diagonal entry of $Y$. Thus, all results in Section 3 extend verbatim, and gradient descent exhibits chaos and a fractal convergence boundary on $\mathcal{W}$ (see details in Appendix G).
A full characterization of the dynamics outside $\mathcal{W}$ requires additional investigations, for instance, whether $\mathcal{W}$ attracts or repels nearby trajectories. However, we point out that, the presence of chaos on $\mathcal{W}$ already suggests that the global dynamics can be unpredictable. For instance, due to the continuity of $\mathrm{GD}_{\eta}$, we expect that initializations in the vicinity of the convergence boundary in $\mathcal{W}$, will inherit the sensitivity to initial conditions, at least during the initial phase before potentially escaping $\mathcal{W}$. This phenomenon is known as transient chaos in dynamical systems (see, e.g., Tél, 1990). Experimentally, we observe that the chaotic phenomena indeed persist for general initializations, and also in deep matrix factorization (Appendix I). A rigorous characterization of gradient descent dynamics in these settings is an intriguing direction that we leave for future work.
4.2 General Mechanism Behind Chaos
We now explain the mechanism giving rise to chaos in scalar factorization. For simplicity, consider the case $d=1$: $L(u,v)=(uv-y)^{2}/2+\lambda(u^{2}+v^{2})/2$ with $(u,v)\in\mathbb{R}^{2}$. In this setting, there exists a set $\mathcal{C}\subset\mathbb{R}^{2}$ such that $\mathrm{GD}_{\eta}$ behaves as a $3$-covering map from $\mathcal{C}$ onto $\mathrm{GD}_{\eta}(\mathcal{C})\supset\mathcal{C}$. Roughly speaking, $\mathrm{GD}_{\eta}$ stretches and folds $\mathcal{C}$ to cover $\mathrm{GD}_{\eta}(\mathcal{C})$ three times. Furthermore, $\mathcal{C}$ contains the convergence boundary $\partial\mathcal{D}_{\eta}$. The boundary satisfies $\mathrm{GD}_{\eta}(\partial\mathcal{D}_{\eta})=\partial\mathcal{D}_{\eta}$, meaning that $\partial\mathcal{D}_{\eta}$ can be stretched and folded three times to cover itself, thereby exhibiting self-similarity. Additionally, due to its folding behavior, $\mathrm{GD}_{\eta}$ is transitive on $\partial\mathcal{D}_{\eta}$, i.e., points on the boundary $\partial\mathcal{D}_{\eta}$ are mixed under the iterations of $\mathrm{GD}_{\eta}$ (via multiple stretches and folds). This mixing leads to the sensitivity to initializations near the boundary. A visualization of these properties are provided in Figure 6.
The key ingredient above is the existence of a set $\mathcal{C}$ in the parameter space on which $\mathrm{GD}_{\eta}$ acts as a covering map. In fact, the described chaotic phenomena arise given such a set $\mathcal{C}$ and any invariant subset $\mathcal{A}\subset\mathcal{C}$ that satisfies $\mathrm{GD}_{\eta}(\mathcal{A})=\mathcal{A}$, a condition typically satisfied by the boundary of a basin of attraction. Then we ask: When does such a set $\mathcal{C}$ exist? The following result partially addresses this for general neural networks with polynomial activation functions.
@@ 4.2 General Mechanism Behind Chaos
Let $f_{\theta}(\cdot)$ be a polynomial neural network of arbitrary depth and width, parametrized by $\theta\in\mathbb{R}^{p}$. Consider a loss function $L(\theta)=\sum_{j=1}^{m}\ell(f_{\theta}(x_{j}),y_{j})$ with polynomial loss $\ell$ and training data $(x_{j},y_{j})_{j=1}^{m}$. For all $\eta>0$ except for at most finitely many values, there exists a measure-zero set $\mathcal{K}_{\eta}\subset\mathbb{R}^{p}$ such that $\mathrm{GD}_{\eta}$ is a covering map on any connected component of $\mathbb{R}^{p}\setminus\mathcal{K}_{\eta}$.
The proof is provided in Appendix G. We remark that, while Proposition 6 suggests a general folding behavior of $\mathrm{GD}_{\eta}$, it does not address whether $\mathrm{GD}_{\eta}(\mathcal{C})$ contains $\mathcal{C}$ and whether $\mathcal{C}$ contains a basin boundary. We leave these intriguing questions for future research.
4.3 Chaos and Fractals in Neural Networks
We show that the chaotic phenomena persist in real-world training settings. We trained a depth-three ReLU network on a $2$-class subset of CIFAR-10 (Krizhevsky et al., 2009). In Figure 4 we consider the mean squared error and report how the step size affects the norm and sharpness of the parameter that is returned at the end of training. For gradient descent without momentum, two distinct regimes of step sizes can be observed: (i) EoS regime: both the final norm and sharpness lie close to a smooth curve, indicating predictability. In particular, the final sharpness is close to $2/\eta$, aligning with Cohen et al. (2021). (ii) Chaotic regime: the norm and sharpness become sensitive to the step size, indicating chaos and unpredictability. In particular, the final sharpness spans almost all values below $2/\eta$, aligning with our Theorem 2. A similar phenomenon appears for gradient descent with Polyak momentum (Polyak, 1964). However, in the small-step-size regime, the final sharpness forms a cluster rather than aligning with $(2+2\beta)/\eta$ as predicted by Cohen et al. (2021). Experiments with cross-entropy loss are provided in Appendix J, where a similar two-regime phenomenon is observed.
Figure 4: GD without and with momentum in training a depth-3 ReLU network on a subset of CIFAR-10 for $5000$ iterations under mean squared error. The initialization is randomly sampled once and then kept fixed across all panels. At large step sizes, the final norm and final sharpness are sensitive to step size. Dashed black lines show the final sharpness predicted by Cohen et al. (2021).
In Figure 5, we show how the parameter initialization of gradient descent affects the final loss and sharpness when using weight decay. For both quantities, we observe fractal structures in the parameter space, indicating that the training outcome is highly sensitive to the initialization. Further experiments in Appendix J show qualitatively similar fractal patterns also appear without weight decay. Experiment details of Figures 4 and 5 are provided in Appendix J.
Figure 5: GD with weight decay in training a depth-3 ReLU network on a subset of CIFAR-10 for $3000$ iterations. Shown is a random two-dimensional slice of the parameter space. Each initial parameter is colored by the value of the final loss and final sharpness, respectively. Left two panels: fractal basins of global minimizers (dark blue), sub-optimal solutions (white), and a region leading to divergence (dark red). Right two panels: the final sharpness is sensitive to the initialization, spanning a wide range of values below $2/\eta$. Gray points are initializations from which the algorithm diverged.
@@ 5 Conclusion
We offered a rigorous characterization of gradient descent with large step sizes in matrix factorization. Our results reveal two striking phenomena: near the convergence boundary, the selection of the minimizer is unpredictable, and adding regularization can induce a fractal convergence boundary that makes the convergence itself unpredictable. As a driver of this complexity, we suggested a covering map structure exhibited by the gradient descent update map on the parameter space.
@@
Although our characterizations substantially expand the state of knowledge in non-convex overparametrized optimization in the particular setting of matrix factorization, further research is needed to rigorously characterize the dynamics of large-step-size gradient descent in other settings, such as general initializations, deep matrix factorization, or neural networks with nonlinear activation functions. We believe the contributed insights can aid in the development of such programs.
@@
Code to reproduce our experiments is made available at https://github.com/shuangliang15/chaos-matrix-factorization.
@@ Acknowledgment
This project has been supported by NSF CCF-2212520 and NSF DMS-2145630. GM was partially supported by DARPA AIQ project HR00112520014, DFG SPP 2298 project 464109215, and BMFTR in DAAD project 57616814 (SECAI). We thank Ryan Tibshirani for insightful discussions.
@@
  • Appendix I: Additional experiments on matrix factorization
  • Appendix J: Additional experiments on real-world data
@@ A.1 Chaotic dynamical systems
A common definition of chaotic dynamical systems is as follows (Devaney and Eckmann, 1987). A dynamical system $F\colon X\to X$, where $X$ is the state space, is chaotic if: (i) it is sensitive to initialization, (ii) it is topological transitive, and (iii) periodic points are dense in $X$. Here, sensitivity to initialization requires that there exists $\delta>0$ such that for any $x\in X$ and any neighborhood of $x$, there exists $N>0$ and $y$ in the neighborhood such that $\operatorname{dist}(F^{N}(x),F^{N}(y))>\delta$. This is weaker than the property we presented in Theorem 1 and Theorem 4: when the converged points of two trajectories are different, the trajectories must differ by a positive difference at some time $N$, but not vice versa. In Proposition 15, we show that the boundary $\partial\mathcal{D}_{\eta}^{\prime}$, as defined in Theorem 1, is invariant under gradient descent. In Proposition 18, we show that when restricted to $\partial\mathcal{D}_{\eta}^{\prime}$, the gradient descent system is semi-conjugate to a one-dimensional system that is precisely Devaney chaotic. However, we show in Proposition 20 that the original gradient descent system is not Devaney chaotic when $d\geq 2$ as it fails to be topological transitive.
A system $F\colon X\to X$ is topological transitive if for any pair of non-empty open sets $U,V$, there exists $N$ such that $F^{N}(U)\cap V\neq\varnothing$. A family of dynamical systems that exhibit transitivity is the family of Axiom-A diffeomorphisms (Smale, 1967). These are dynamical systems where the set of non-wandering points is hyperbolic and is equal to the closure of the set of periodic points. A closed set $\Lambda$ is hyperbolic if it is forward invariant and at each point $x\in\Lambda$ the tangent space of the ambient space splits as a direct sum of stable and unstable subspaces. It is known that an Axiom-A diffeomorphism is always transitive on each of its basic sets (Bowen et al., 2008, Chapter 3). However, gradient update maps typically are not expected to satisfy this definition as they in general are not global diffeomorphisms.
Another definition of chaotic dynamical system is by Li and Yorke (1975). They considered a dynamical system $F\colon X\to X$ with $X\subset\mathbb{R}$ being an interval chaotic if $F$ has a periodic orbit with period three. They showed that if such a periodic orbit exists, then (i) $F$ has periodic orbit with any period; (ii) there exists an uncountable set $S\subset J$ such that, for every $p,q\in S$ with $p\neq q$,
@@
In Proposition 18, we showed that the restricted system $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}}$ is semi-conjugate to a one-dimensional system that is Li-Yorke chaotic. In general, Devaney chaos and Li-Yorke chaos do not imply each other. For a detailed comparison between different notions of chaotic dynamical systems, see the work of Elaydi (2007).
@@ A.2 Fractal basin boundary
The fractal convergence boundary studied in this work falls into a more general notion, called a fractal basin boundary. The seminal classification given by McDonald et al. (1985) divides fractal basin boundaries into three categories: quasicircles, locally connected but not quasicircles, and locally disconnected. The most regular type, quasicircle, is typical in the Julia sets of complex analytic maps. However, as noted by McDonald et al. (1985), properties of complex analytic maps do not generalize to real maps, and hence quasicircle is uncommon in real systems. We observe that the convergence boundary in our study falls in the second category, locally connected but not quasicircles. This type of boundary has been observed in several planar maps, i.e., dynamical systems defined on regions of $\mathbb{R}^{2}$. A well-known example is the following system:
@@
The basin boundary of this system is precisely the Weierstrass curve. McDonald et al. (1985) argued that a typical characteristic of this type of boundaries is the local stratification structure, which also appears in the convergence region in our case (see left panel in Figure 1). To our knowledge, however, all examples of locally connected boundaries appearing in the literature, including those presented by McDonald et al. (1985); Hunt et al. (1999); Rosa Jr and Ott (1999), are bounded, whereas the convergence region in our case is shown to be unbounded. Classical approaches do not apply to our study, as most of those theoretical studies are case-specific. The last category has the most complicated structure and, as noted by Aguirre et al. (2009), turns out to appear more commonly in physical systems. Boundaries in this category typically exhibit a Cantor set structure. Examples include the famous Hénon map and the horseshoe map. For a recent review of the fractal boundaries, we refer readers to Aguirre et al. (2009).
Appendix B Measure of chaos in dynamical systems